OpenStudy (anonymous):

How do you find the equation of a line that goes through 2 points?

OpenStudy (anonymous):

first find the slope using \[m=\frac{y_2-y_1}{x_2-x_1}\] then use the almighty point slope formula via \[y=y_1=m(x-x_1)\]

OpenStudy (anonymous):

hello polpak!

OpenStudy (anonymous):

First find the slope of the line using the slope formula: \[Slope = \frac{y_2-y_1}{x_2-x_1}\] Where (x1, y1) and (x2,y2) are your two points. Then take that slope and plug it in for m along with the x,y values from one of your points into the point-slope formula: \[y - y_1 = m(x-x_1)\]

OpenStudy (anonymous):

Hi Satellite